OpenStudy (xhawaiiangirlx):

How many solutions does the equation 4x + 8 = −2 − 2 + 5x have? Two One None Infinitely many @Frozen15

OpenStudy (priyar):

do u know how to solve this eq?

OpenStudy (xhawaiiangirlx):

no

OpenStudy (xhawaiiangirlx):

I suck at math

OpenStudy (priyar):

try to write in the form of x=.. by bringing all the constants(numbers) to one side..don't forget to change the sign when u do so

OpenStudy (priyar):

u got it?

OpenStudy (xhawaiiangirlx):

x=4x + 8 = −2 − 2 + 5x

OpenStudy (priyar):

an equation can't have 2 '=' signs!

OpenStudy (priyar):

-2-2=-4 take it to lhs... so 8+4=12

OpenStudy (priyar):

4x+12=5x now take 4x to rhs ...so sign changes it becomes -4x...so 5x-4x=x...

OpenStudy (priyar):

now tell me what will be the answer..?

OpenStudy (frozen15):

lol these kind of questions always turn out to be So many equations , plus this looks very familiar so D is correct for sure just to help

OpenStudy (priyar):

why D ? @Frozen15

OpenStudy (priyar):

Did you get the answer? @XhawaiiangirlX